Tge hypotenuse of a right triangle measures 10cm. One leg is 2 cm shorter than the other. What are the lengths of the legs? Use the quadratic formula.
Answer by mananth(16949)

Let one leg be x
the other leg = x-2
x^2 + (x-2)^2= 10^2
x^2+x^2-4x+4=100
2x^2-4x-96=0 divide equation by 2
x^2-2x-48=0
x^2-8x+6x-48=0
x(x-8)+6(x-8)=0
(x-8)(x+6)=0
x=8 0r -6
So x is 8 which is a positive value